OpenStudy (anonymous):

find the value for the given ratio 2x/5=14/30

OpenStudy (anonymous):

\[\Large \begin{array}{l} \frac{{2x}}{5} = \frac{{14}}{{30}}\\ \frac{{12x}}{{30}} = \frac{{14}}{{30}}\\ x = \frac{{14}}{{12}} = \frac{7}{6} \end{array}\]

OpenStudy (anonymous):

where did you get the 12?

OpenStudy (anonymous):

I'm finding a common denominator for the two fractions. Note that 5 goes into 30 6 times, and 6*2 = 12. Whatever you do to the bottom, you have to do to the top.

OpenStudy (anonymous):

multiply by 5, divide by 2

OpenStudy (anonymous):

what is this "large" motif?

OpenStudy (anonymous):

so seven is the top numerator right?

OpenStudy (anonymous):

@satellite73: it makes it bigger and a bit easier to read, sometimes MathJax doesn't render the equations nicely when they're small (I mean they're proper, but numbers appear to run into each other and are sometimes hard to read).